A child whose mass is 50 kg sits on a rectangular swimming raft 1 meter long and 1 meter wide that is floating in freshwater. By how much does the raft rise after he dives from it?
1
Expert's answer
2018-07-20T15:37:08-0400
The weight of the child is compensated by the Archimedes’ force. Thus
mg=ρgV,
where g – acceleration due to gravity, ρ=1000 kg/m^3 – density of water, V – immersed volume of the raft.
Hence V=m/ρ. Since the area of the raft is A=1 m^2, the raft rises by V/A=m/ρA=5 cm after the child dives from it.
Answer: 5 cm.